3. Bricks are delivered to a work site and stacked in rows and columns, forming a rectangular prism. The length of the prism is 1 foot greater than its width, and its height is 2 feet less than its width. Show how to find the dimensions of the prism formed by the bricks, given that its volume is 40 cubic feet.
Answer by CubeyThePenguin(3113)

w = width

w(w + 1)(w - 2) = 40

factors of 40: 1, 2, 4, 5, 8, 10, 20, 40

Pick 3 factors of 40 that multiply to 40: 2, 4, 5.

width = w = 4 ft
length = w + 1 = 5 ft
height = w - 2 = 2 ft